Determine the clearance at the lowest point of the conductor with respect to the ground from a hilly terrain which has a gradient of 1 in 20 between two towers of 36m in height and having a span length of 300m. An overhead line is supported by these towers wherein the lowest conductor is fixed 6m below the top of each tower. The conductor weigh is 1.5kgf per metre and the tension is 2200kgf per metre. Assume a parabolic configuration.
Answer #1

Given Gradient 1:20: H=S-si H f S₂ ना Si A G o D F the conductors are supported between Towers AD and Be over hill side hevinL Vertical distance between Towers is h = EC = DE Sino = 3oox 1/20 = 15m. Horizontal distance between two towers is =/300²_15X, + X2 - 300 X, - 300 – 223033 x = 76.66 m Sag f₂ = W X 2 1X 223.33 AT = 11:33 m 2 X 2200 Bc = BET Ec = 30+15=45m. Clearonce